# Scalable magnetic resonance fingerprinting: Incremental inference of high dimensional elliptical mixtures from large data volumes

### v1

@@ 5.2 Cluster globally, reduce locally
The proposed dimension reduction and reconstruction, using the HD-ED model, generalizes to GSM, PPCA (Tipping and Bishop, 1999c) and robust PPCA (Archambeau, Delannay and Verleysen, 2006), the former using Gaussian and the latter Student distributions. Both have been extended to account for potential heterogeneity in data, considering mixtures, with the MPPCA model (Tipping and Bishop, 1999d) and its Student-based robust version (Archambeau, Delannay and Verleysen, 2008). In these models, all clusters are assumed to live in subspaces of the same dimension $d$. In the Gaussian case, an extension, allowing varying reduced dimensions $d_{k}$ across clusters, have been proposed by Bouveyron, Girard and Schmid (2007) with their high dimensional Gaussian mixture model (HD-GMM). The proposed HD-GMM parameterization allows to handle high dimensional data in a computationally efficient way. However, it does not provide an actual lower dimensional representation of the data. While such a reduced-dimensional representation may often not be needed, it may be crucial to deal with hardware or software limitations. Originally, HD-GMM have not been designed for dimension reduction or compression but rather for clustering and density estimation in high-dimensional heterogeneous settings. Our HD-MED model uses the same efficient decomposition of the covariance matrix but generalizes it to the scale matrix of a GSM distribution. In addition, we describe how it can be further exploited as a dimension reduction technique. As finite mixture models, HD-MED can be used for clustering data into $K$ clusters. For any possible observation $\mathbf{y}$, a HD-MED model provides a probability $r_{k}(\mathbf{y})$ that $\mathbf{y}$ is assigned to cluster $k$ for each $k\in[1\!:\!K]$. Using (9), a reduced-dimension representation $\widehat{\mathbf{y}}_{k}$ of $\mathbf{y}$, for each of the $K$ different subspaces, is denoted by $\widehat{\mathbf{y}}_{k}=Q_{k}(\mathbf{y})$ and given by,
@@
while its reconstruction $\check{\mathbf{y}}_{k}$ in the original space is given by
$\check{\mathbf{y}}_{k}={\mathb{V}}_{k}\widehat{\mathbf{y}}_{k}+{\boldsymbol{\mu}}_{k}\;.$
In practice, it is reasonable to use as a reduced-dimension representation of $\mathbf{y}$ only the one corresponding to the most probable cluster $k$, i.e. with the highest $r_{k}(\mathbf{y})$. In this setting, HD-MED acts as a divide-and-conquer paradigm by initially clustering the data into $K$ clusters and then performing cluster-specific data reduction. The divide step allows a much more effective reduction than if a single subspace was considered, while in the conquer step, little information is lost, as for any new observation $\mathbf{y}$, cluster assignment probabilities $r_{k}(\mathbf{y})$ can be straightforwardly computed to decide on the best reduced representation to be used. However, for subsequent processing, it is important to keep track of clustering information for each observation. The reduced representations cannot be pooled back altogether, as they are likely to become impossible to distinguish across clusters. Also, as summarized in Algorithm 1 and illustrated in Section 7, each reduced cluster may have to be processed separately but this additional cost is negligible compared to the hardware and software gain of a more efficient representation.
In practice, the Expectation-Maximization (EM) algorithm (Dempster, Laird and Rubin (1977)), that iteratively computes and maximizes the conditional expectation of the complete-data log-likelihood, is commonly used to infer the parameters of finite mixture models. The number of clusters $K$, and their respective inner dimension $d_{k}$ are hyper-parameters that need to be tuned prior to the EM steps. As a simple solution, we use the Bayesian Information Criterion (BIC) to tune $K$ and the $d_{k}$’s at the same time. In contrast to MPPCA solutions which assume the same subspace dimension $d$ for all clusters, the possibility to handle different $d_{k}$’s and to allow non Gaussian cluster shapes is important for the target applications involving datasets that are very large. Using different dimensions across clusters is likely to yield a more efficient reduced representation of the data as illustrated in Section 7.
However, the standard, or batch EM algorithm needs all the dataset to be loaded in a fast access memory (e.g the RAM) which is often limited as this kind of memory is expensive. In the case of large-scale dataset, the RAM is often overloaded and supported by a slow-memory which makes the iterations of the EM very slow. Batch sizes are then limited by resource constraints, so that very large data sets need either to be downsampled or to be handled in an incremental manner. Incremental versions of EM exist and can be adapted to our setting. In section 6, we provide a way to deal with this large-scale case by using an online version of the EM algorithm.
@@ 6.1 Online EM, main assumptions
When the data volume is too large the EM algorithm becomes slow because of multiple data transfer between the RAM and the store of the computer. A way to handle large volumes is to use online learning. Online learning refers to procedures able to deal with data acquired sequentially. Online variants of EM, among others, are described in Cappé and Moulines (2009); Maire, Moulines and Lefebvre (2017); Karimi et al. (2019a, b); Fort, Moulines and Wai (2020); Kuhn, Matias and Rebafka (2020); Nguyen, Forbes and McLachlan (2020). As an archetype of such algorithms, we consider the online EM of Cappé and Moulines (2009) which belongs to the family of stochastic approximation algorithms (Borkar (2009)). This algorithm has been well theoretically studied and extended. However, it is designed only for distributions that admit a data augmentation scheme, or a latent variable formulation, yielding a complete likelihood of the exponential family form, see (12) below. This case is already very broad, including e.g. Gaussian, gamma, Student distributions and mixtures of those. We recall the main assumptions required and the online EM iteration, based on a latent variable formulation.
Assume $\left(\mathbf{Y}_{i}\right)_{i=1}^{N}$ is a sequence of $N$ i.i.d. replicates of a random variable $\mathbf{Y}\in{\cal Y}\subset\mathbb{R}^{M}$, observed one at a time. Extension to successive mini-batches of observations is straightforward (Nguyen, Forbes and McLachlan (2020)). In addition, $\mathbf{Y}$ is assumed to be the visible part of $\left(\mathbf{Y},\mathbf{Z}\right)$, where $\mathbf{Z}\in\mathbb{R}^{l}$ is a latent variable, e.g. the unknown component label in a mixture model or a mixing weight in a GSM formulation, and $l\in\mathbb{Z}_{+}$. For $i\in[1:N]$ then, each $\mathbf{Y}_{i}$ is the visible part of $\left(\mathbf{Y}_{i},\mathbf{Z}_{i}\right)$. Suppose $\mathbf{Y}$ arises from some data generating process (DGP) characterised by a probability density function $f\left(\mathbf{y};\boldsymbol{\Theta}_{0}\right)$, with unknown parameters $\boldsymbol{\Theta}_{0}\in\mathbb{T}\subseteq\mathbb{R}^{p}$, for $p\in\mathbb{Z}_{+}$.
Using the sequence $\left(\mathbf{Y}_{i}\right)_{i=1}^{N}$, the method of Cappé and Moulines (2009) allows to sequentially estimate $\boldsymbol{\Theta}_{0}$ provided the following assumptions are met:
@@
Let $\left(\gamma_{i}\right)_{i=1}^{N}$ be a sequence of learning rates in $\left(0,1\right)$ and let $\boldsymbol{\Theta}^{\left(0\right)}\in\mathbb{T}$ be an initial estimate of $\boldsymbol{\Theta}_{0}$. For each $i\in[1:N]$, the online EM of Cappé and Moulines (2009) proceeds by computing
@@
and
$\boldsymbol{\Theta}^{\left(i\right)}=\bar{\boldsymbol{\Theta}}(\mathbf{s}^{\left(i\right)})\text{,}$ (16)
where $\mathbf{s}^{\left(0\right)}=\bar{\mathbf{s}}(\mathbf{y}_{1};\boldsymbol{\Theta}^{\left(0\right)})$. It is shown in Theorem 1 of Cappé and Moulines (2009) that when $N$ tends to infinity, the sequence $(\boldsymbol{\Theta}^{\left(i\right)})_{i=1:N}$ of estimators of $\boldsymbol{\Theta}_{0}$ satisfies a convergence result to stationary points of the likelihood (cf. Cappé and Moulines (2009) for a more precise statement).
@@ 6.2 Online EM for HD-ED
In this subsection, we derive the online EM algorithm for HD-ED. The extension to mixtures of those (HD-MED) is straightforward and is detailed in Nguyen and Forbes (2022) or in the supplementary Section 1. The weight $W$ distribution $f$ in the GSM formulation (3) is assumed to belong to the exponential family. This case may seem restrictive but it encompasses a number of ED as the Gaussian, Student, Normal Inverse Gamma with no skewness etc. distributions.
@@
where $\operatorname{vec}$ denotes the vectorization operator (Schott (2016)).
@@
The online EM algorithm (OEM) consists, as the batch EM, of two steps, the first one is the computation of the sufficient statistics (13), and the second one is the maximization of the likelihood (14). For the first step, we need to compute $\overline{{\mathb{s}}}\left(\mathbf{y};{\boldsymbol{\Theta}}\right)=\mathbb{E}[{\mathb{s}}(\mathbf{Y},\mathbf{Z})|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}]$. This quantity requires to compute the following expectations $\mathbb{E}[W|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}]$, and $\mathbb{E}[s_{w}(W)|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}]$.
@@
Let $\mathbf{Y}\sim\mathcal{HE}_{Md}({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}})$, ${\boldsymbol{\Theta}}=\left({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}}\right)$, and $W$ the mixing variable $W\sim f_{{\boldsymbol{\theta}}}$. Then $\mathbf{Y}$ has density (2) with generator $g$ and Mahalanobis distance $u=(\mathbf{y}-{\boldsymbol{\mu}})^{T}{\mathb{D}}{\mathb{A}}^{-1}{\mathb{D}}^{T}(\mathbf{y}-{\boldsymbol{\mu}})$ defined as in (7). It follows that
@@
According to equations (2) and (3), we have
$\mathbb{E}\left[W|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}\right]$ $=\frac{1}{(2\pi)^{M/2}g(u)}\int\limits_{\mathbb{R}^{+}}w\;f_{{\boldsymbol{\theta}}}(w)\exp(-\frac{w}{2}u)\mathop{}\!\mathrm{d}w$
$=-\frac{2}{(2\pi)^{M/2}}\frac{g^{\prime}(u)}{g(u)},$
which proves equation (19). ∎
@@
with ${\mathb{s}}_{1}$ a $M$-dimensional vector, ${\mathb{S}}_{2}$ a $M\times M$ matrix, and $s_{3}$, $s_{4}$, $s_{5}$ three scalar values. Parameters are updated by maximizing $Q$ with respect to ${\boldsymbol{\Theta}}$. $\bar{{\boldsymbol{\Theta}}}({\mathb{s}})$ is defined as the root of the first-order condition
${\mathb{J}}_{{\boldsymbol{\phi}}}({\boldsymbol{\Theta}}){\mathb{s}}-\frac{\partial\psi}{\partial{\boldsymbol{\Theta}}}({\boldsymbol{\Theta}})=\mathbf{0},$ (21)
where ${\mathb{J}}_{{\boldsymbol{\phi}}}({\boldsymbol{\Theta}})=\frac{\partial{\boldsymbol{\phi}}}{\partial{\boldsymbol{\Theta}}}$ is the Jacobian of $\phi$, with respect to ${\boldsymbol{\Theta}}$. Computing gradients leads to $\bar{{\boldsymbol{\Theta}}}({\mathb{s}})=(\overline{{\boldsymbol{\mu}}}({\mathb{s}}),\overline{\widetilde{{\mathb{D}}}^{*}}({\mathb{s}}),\overline{{\mathb{A}}}({\mathb{s}}),\overline{{\boldsymbol{\theta}}}({\mathb{s}})$, where $\overline{{\boldsymbol{\mu}}}({\mathb{s}})$, and $\overline{{\mathb{A}}}({\mathb{s}})$ are closed form, and $\overline{\widetilde{{\mathb{D}}}^{*}}({\mathb{s}})$ can be found using Riemannian optimization.
@@
We proceed in a ECM-like procedure (Meng and Rubin (1993)) by optimizing each parameters separately and incorporating them during the optimization of each other parameters. $({\boldsymbol{\mu}},{\mathb{A}})$ can be optimized easily using (21) and computing the gradients gives
@@
The maximisation in $\widetilde{{\mathb{D}}}^{*}$ has to take into account that $\widetilde{{\mathb{D}}}^{*}\in St(M,d)$, the Stiefel manifold of the $M\times d$ matrices. Plugin-in the expressions of ${\boldsymbol{\mu}}$ and ${\mathb{A}}$ above and omitting parts that depend on ${\boldsymbol{\theta}}$ we have,
@@
We compute the gradients and use (21) for ${\boldsymbol{\mu}}$ and ${\mathb{A}}$, however for the vector $\widetilde{{\mathb{D}}}^{*}$ we only plug-in the optimized values $\overline{{\boldsymbol{\mu}}}$, and $\overline{{\mathb{A}}}$ in (17) and solve
$\mathop{\mathrm{arg\,max}}\limits_{\widetilde{{\mathb{D}}}^{*}\in St(M,d)}{\mathb{s}}^{T}{\boldsymbol{\phi}}\left(\overline{{\boldsymbol{\mu}}},\overline{{\mathb{A}}},\widetilde{{\mathb{D}}}^{*}\right)-\psi\left(\overline{{\boldsymbol{\mu}}},\overline{{\mathb{A}}},\widetilde{{\mathb{D}}}^{*}\right).$ (27)
Algorithm 1 Divide & Conquer high dimensional matching for MRF reconstruction
@@
1:Online HD-MED inference from $\{\mathbf{y}_{i}\}_{i=1:N}$: $K$ clusters, $d_{k}<M$ for $k\in 1:K$$\Longrightarrow$ cluster assignment probabilities and cluster-wise projections $({\mathb{r}},{\mathb{Q}})=\{r_{k}(\cdot),Q_{k}(\cdot)\}_{k=1:K}$
@@
2:Matching: For $k\in 1:K$, for $j\in\tilde{I}_{k}$, determine $i(\tilde{\mathbf{y}}_{j})=\arg\min\limits_{i\in I_{k}}d(Q_{k}(\widetilde{\mathbf{y}}_{j}),\widehat{\mathbf{y}}_{i})$ and set $\widetilde{\mathbf{t}}_{j}=\mathbf{t}_{i(\tilde{\mathbf{y}}_{j})}$

### v2

@@ 5.2 Cluster globally, reduce locally
The proposed dimension reduction and reconstruction, using the HD-ED model, generalizes to GSM, PPCA (tipping1999) and robust PPCA (archambeau2006), the former using Gaussian and the latter Student distributions. Both have been extended to account for potential heterogeneity in data, considering mixtures, with the MPPCA model (tipping1999mix) and its Student-based robust version (archambeau2008). In these models, all clusters are assumed to live in subspaces of the same dimension $d$. In the Gaussian case, an extension, allowing varying reduced dimensions $d_{k}$ across clusters, have been proposed by bouveyron2007 with their high-dimensional Gaussian mixture model (HD-GMM). The proposed HD-GMM parameterization allows to handle high-dimensional data in a computationally efficient way. However, it does not provide an actual lower dimensional representation of the data. While such a reduced-dimensional representation may often not be needed, it may be crucial to deal with hardware or software limitations. Originally, HD-GMM have not been designed for dimension reduction or compression but rather for clustering and density estimation in high-dimensional heterogeneous settings. Our HD-MED model generalizes the covariance matrix decomposition to the scale matrix of a GSM distribution. In addition, we describe how it can be further exploited as a dimension reduction technique. As mixture models, HD-MED can be used for clustering data into $K$ clusters. For any possible observation $\mathbf{y}$, a HD-MED model provides a probability $r_{k}(\mathbf{y})$ that $\mathbf{y}$ is assigned to cluster $k$ for each k=1:Kk=1\!:\!K. Using (9), a reduced-dimension representation $\widehat{\mathbf{y}}_{k}$ of $\mathbf{y}$, for each of the $K$ different subspaces, is denoted by $\widehat{\mathbf{y}}_{k}=Q_{k}(\mathbf{y})$ and given by,
@@
while its reconstruction $\check{\mathbf{y}}_{k}$ in the original space is given by $\check{\mathbf{y}}_{k}={\mathb{V}}_{k}\widehat{\mathbf{y}}_{k}+{\boldsymbol{\mu}}_{k}\;.$ In practice, it is reasonable to use as a reduced-dimension representation of $\mathbf{y}$ only the one corresponding to the most probable cluster $k$, i.e. with the highest $r_{k}(\mathbf{y})$. In this setting, HD-MED acts as a divide-and-conquer paradigm by initially clustering the data into $K$ clusters and then performing cluster-specific data reduction. The divide step allows a much more effective reduction than if a single subspace was considered, while in the conquer step, little information is lost, as for any new observation $\mathbf{y}$, cluster assignment probabilities $r_{k}(\mathbf{y})$ can be straightforwardly computed to decide on the best reduced representation to be used. However, it is important to keep track of clustering information for each observation. The reduced representations cannot be pooled back altogether, as they are likely to become impossible to distinguish across clusters. Also, as summarized in Algorithm 1 and illustrated in Section 7, each reduced cluster may have to be processed separately but this additional cost is negligible compared to the hardware and software gain of a more efficient representation.
The Expectation-Maximization (EM) algorithm (dempster1977) that iteratively maximizes the conditional expectation of the complete-data log-likelihood, is commonly used to estimate finite mixture models. Gradient algorithms can also be used but have not been yet very popular, see a more detailed discussion in Appendix Section 2. The number of clusters $K$, and their respective inner dimension $d_{k}$ are hyperparameters that need to be tuned prior to the EM steps. As a simple solution, we use the Bayesian Information Criterion (BIC) to tune $K$ and the $d_{k}$’s at the same time. In contrast to MPPCA solutions which assume the same subspace dimension $d$ for all clusters, the possibility to handle different $d_{k}$’s and to allow non Gaussian cluster shapes is important for the target applications involving datasets that are very large. Using different dimensions across clusters is likely to yield a more efficient reduced representation of the data as illustrated in Section 7.
The standard, or batch EM algorithm needs all the dataset to be loaded in a fast access memory (e.g the RAM) which is often limited as this kind of memory is expensive. In the case of large-scale dataset, the RAM is often overloaded and supported by a slow-memory which makes EM iterations very slow. Batch sizes are then limited by resource constraints, so that very large data sets need either to be downsampled or to be handled in an incremental manner. Incremental versions of EM exist and can be adapted to our setting. In section 6, we provide a way to deal with this large-scale case by using an online version of EM.
@@ 6.1 Online EM, main assumptions
When the data volume is too large, EM becomes slow because of multiple data transfer between the RAM and the computer storage. A way to handle large volumes is to use online learning. Online learning refers to procedures dealing with data acquired sequentially. Online variants of EM, among others, are described in Cappe2009; Maire:2017aa; Karimi:2019ab; Karimi:2019aa; Fort:2020aa; Kuhn:2020aa; Nguyen:2020ac; Dieuleveut2023. As an archetype of such algorithms, we consider the online EM of Cappe2009 which belongs to the family of stochastic approximation algorithms (borkar2009; Dieuleveut2023). This algorithm has been theoretically studied and extended. However, it is designed only for distributions that admit a data augmentation scheme yielding a complete likelihood of the exponential family form, see (12) below. This case is already very broad, including e.g. Gaussian, gamma, Student distributions and mixtures of those. We recall the main assumptions required and the online EM iteration, based on a latent variable formulation.
Assume $\left(\mathbf{Y}_{i}\right)_{i=1}^{N}$ is a sequence of $N$ i.i.d. replicates of a random variable $\mathbf{Y}\in{\cal Y}\subset\mathbb{R}^{M}$, observed one at a time. Extension to successive mini-batches is straightforward (Nguyen:2020ac). In addition, $\mathbf{Y}$ is assumed to be the visible part of $\left(\mathbf{Y},\mathbf{Z}\right)$, where $\mathbf{Z}\in\mathbb{R}^{l}$ is a latent variable, e.g. the unknown component label in a mixture or a mixing weight in a GSM formulation, and $l\in\mathbb{Z}_{+}$. For $i\in[1:N]$ then, each $\mathbf{Y}_{i}$ is the visible part of $\left(\mathbf{Y}_{i},\mathbf{Z}_{i}\right)$. Suppose $\mathbf{Y}$ arises from some data generating process (DGP) characterised by a probability density function $f\left(\mathbf{y};\boldsymbol{\Theta}_{0}\right)$, with unknown parameters $\boldsymbol{\Theta}_{0}\in\mathbb{T}\subseteq\mathbb{R}^{p}$, for $p\in\mathbb{Z}_{+}$.
Using the sequence $\left(\mathbf{Y}_{i}\right)_{i=1}^{N}$, the method of Cappe2009 allows to sequentially estimate $\boldsymbol{\Theta}_{0}$ provided the following assumptions are met:
@@
Let $\left(\gamma_{i}\right)_{i=1}^{N}$ be a sequence of learning rates in $\left(0,1\right)$ and $\boldsymbol{\Theta}^{\left(0\right)}\in\mathbb{T}$ an initial estimate of $\boldsymbol{\Theta}_{0}$. For each $i\in[1:N]$, the online EM of Cappe2009 proceeds by computing
@@
$\mbox{and}\qquad\boldsymbol{\Theta}^{\left(i\right)}$ $=\bar{\boldsymbol{\Theta}}(\mathbf{s}^{\left(i\right)})\text{,}$ (16)
where $\mathbf{s}^{\left(0\right)}=\bar{\mathbf{s}}(\mathbf{y}_{1};\boldsymbol{\Theta}^{\left(0\right)})$. Theorem 1 of Cappe2009 shows that the sequence $(\boldsymbol{\Theta}^{\left(i\right)})_{i=1:N}$ of estimators of $\boldsymbol{\Theta}_{0}$ satisfies a convergence result to stationary points of the likelihood (cf. Cappe2009 for a more precise statement).
@@ 6.2 Online EM for HD-ED
We now derive the online EM for HD-ED. The extension to HD-MED mixtures is straightforward and is detailed in NguyenForbes2021 or in the supplementary Section 3. The weight distribution $f$ in the GSM formulation (3) is assumed to belong to the exponential family. This case may seem restrictive but it encompasses a number of ED as the Gaussian, Student, Normal Inverse Gamma with no skewness etc. distributions.
@@
where $\operatorname{vec}$ denotes the vectorization operator (Schott2016).
@@
The online EM algorithm (OEM) consists, as the batch EM, of two steps, the computation of the sufficient statistics (13) and the maximization of the likelihood (14). For the first step, we need to compute $\overline{{\mathb{s}}}\left(\mathbf{y};{\boldsymbol{\Theta}}\right)=\mathbb{E}[{\mathb{s}}(\mathbf{Y},\mathbf{Z})|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}]$. This quantity requires to compute the following expectations $\mathbb{E}[W|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}]$, and $\mathbb{E}[s_{w}(W)|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}]$.
@@
Let $\mathbf{Y}\sim\mathcal{HE}_{Md}({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}})$, ${\boldsymbol{\Theta}}=\left({\boldsymbol{\mu}},\widetilde{{\mathb{D}}}^{*},{\mathb{a}},b,{\boldsymbol{\theta}}\right)$, and $W$ the mixing variable $W\sim f_{{\boldsymbol{\theta}}}$. Then $\mathbf{Y}$ has density (2) with generator $g$ and Mahalanobis distance $u=(\mathbf{y}-{\boldsymbol{\mu}})^{T}{\mathb{D}}{\mathb{A}}^{-1}{\mathb{D}}^{T}(\mathbf{y}-{\boldsymbol{\mu}})$ defined as in (7). With $g^{\prime}$ denoting the derivative of $g$, it follows that
@@
Equation (19) results from equations (2) and (3), which lead to
$\mathbb{E}\left[W|\mathbf{Y}=\mathbf{y};{\boldsymbol{\Theta}}\right]=\frac{1}{(2\pi)^{M/2}g(u)}\int\limits_{\mathbb{R}^{+}}w\;f_{{\boldsymbol{\theta}}}(w)\exp(-\frac{w}{2}u)\mathop{}\!\mathrm{d}w=-\frac{2}{(2\pi)^{M/2}}\frac{g^{\prime}(u)}{g(u)}.$
@@
with ${\mathb{s}}_{1}$ an $M$-dimensional vector, ${\mathb{S}}_{2}$ an $M\times M$ matrix, and $s_{3}$, $s_{4}$, $s_{5}$ three scalar values. Parameters are updated by maximizing $Q$ with respect to $\boldsymbol{\Theta}$}. $\bar{{\boldsymbol{\Theta}}}({\mathb{s}})$ is defined as the root of the first-order condition involving ${\mathb{J}}_{{\boldsymbol{\phi}}}({\boldsymbol{\Theta}})=\frac{\partial{\boldsymbol{\phi}}}{\partial{\boldsymbol{\Theta}}}$ the Jacobian of $\phi$,
${\mathb{J}}_{{\boldsymbol{\phi}}}({\boldsymbol{\Theta}}){\mathb{s}}-\frac{\partial\psi}{\partial{\boldsymbol{\Theta}}}({\boldsymbol{\Theta}})=\mathbf{0}.$ (21)
Computing gradients leads to $\bar{{\boldsymbol{\Theta}}}({\mathb{s}})=(\overline{{\boldsymbol{\mu}}}({\mathb{s}}),\overline{\widetilde{{\mathb{D}}}^{*}}({\mathb{s}}),\overline{{\mathb{A}}}({\mathb{s}}),\overline{{\boldsymbol{\theta}}}({\mathb{s}})$, where $\overline{{\boldsymbol{\mu}}}({\mathb{s}})$, and $\overline{{\mathb{A}}}({\mathb{s}})$ are closed form, and $\overline{\widetilde{{\mathb{D}}}^{*}}({\mathb{s}})$ can be found using Riemannian optimization.
@@
We use an ECM-like procedure (meng1993) by optimizing parameters separately and incorporating them during the optimization of each other parameter. $({\boldsymbol{\mu}},{\mathb{A}})$ are easily optimized using (21) and computing the gradients gives
@@
Maximizing over $\widetilde{{\mathb{D}}}^{*}$ has to take into account that $\widetilde{{\mathb{D}}}^{*}\in St(M,d)$, the Stiefel manifold of the $M\times d$ matrices. Using the expressions of $\boldsymbol{\mu}$} and $\mathb{A}$}, omitting parts on $\boldsymbol{\theta}$} we have,
@@
We compute the gradients and use (21) for $\boldsymbol{\mu}$} and $\mathb{A}$}, however for the $\widetilde{{\mathb{D}}}^{*}$ we only plug-in the optimized values $\overline{{\boldsymbol{\mu}}}$, and $\overline{{\mathb{A}}}$ in (17) and solve
$\mathop{\mathrm{arg\,max}}\limits_{\widetilde{{\mathb{D}}}^{*}\in St(M,d)}{\mathb{s}}^{T}{\boldsymbol{\phi}}\left(\overline{{\boldsymbol{\mu}}},\widetilde{{\mathb{D}}}^{*},\overline{{\mathb{A}}},\bar{{\boldsymbol{\theta}}}\right)-\psi\left(\overline{{\boldsymbol{\mu}}},\widetilde{{\mathb{D}}}^{*},\overline{{\mathb{A}}},\bar{{\boldsymbol{\theta}}}\right),$ (27)
which does not depend on the value of $\bar{{\boldsymbol{\theta}}}$. ∎
In practice (25) is solved using a Riemannian optimization framework in the setting where $M>>d$ (wen2013).
Algorithm 1 Divide & Conquer high-dimensional matching for MRF reconstruction
@@
1:Online HD-MED inference from $\{\mathbf{y}_{i}\}_{i=1:N}$: $K$ clusters, $d_{k}<M$ for $k=1:K$ $\quad\Longrightarrow$ cluster assignment probabilities and cluster-wise projections $({\mathb{r}},{\mathb{Q}})=\{r_{k}(\cdot),Q_{k}(\cdot)\}_{k=1:K}$
@@
2:Matching: For $k=1:K$, for $j\in\tilde{I}_{k}$, determine $i(\tilde{\mathbf{y}}_{j})=\arg\min\limits_{i\in I_{k}}d(Q_{k}(\widetilde{\mathbf{y}}_{j}),\widehat{\mathbf{y}}_{i})$ and set $\widetilde{\mathbf{t}}_{j}=\mathbf{t}_{i(\tilde{\mathbf{y}}_{j})}$